Caspian Sea Drinks is considering the purchase of a new water filtration system produced by Rube Goldberg Machines. This new equipment, the RGM-7000, will allow Caspian Sea Drinks to expand production. It will cost $14.00 million fully installed and will be fully depreciated over a 20 year life, then removed for no cost. The RGM-7000 will result in additional revenues of $3.04 million per year and increased operating costs of $507,343.00 per year. Caspian Sea Drinks' marginal tax rate is 25.00%. If Caspian Sea Drinks uses a 10.00% discount rate, then the net present value of the RGM-7000 is _____.

Answer format: Currency: Round to: 2 decimal places.

Assume a par value of $1,000. Caspian Sea plans to issue a 17.00 year, semi-annual pay bond that has a coupon rate of 8.11%. If the yield to maturity for the bond is 7.70%, what will the price of the bond be?

Answer format: Currency: Round to: 2 decimal places.

Assume a par value of $1,000. Caspian Sea plans to issue a 10.00 year, semi-annual pay bond that has a coupon rate of 7.81%. If the yield to maturity for the bond is 8.24%, what will the price of the bond be?

Answer format: Currency: Round to: 2 decimal places

I would really appreciate the help! :) would be great if you could show all steps (not excel) for better understanding~ Thank You.
Answer #1

Answer 1:

The net present value of the RGM-7000 is = $3,661,326.21

Working:

Year 0:

Initial investment = $14 million = $14,000,000

Year 0 to Year 20:

Annual depreciation = (Installed cost - salvage value) / Useful life = (14000000 - 0) / 20 = $700,000

Annual cash flows = (additional revenue - additional operating cost) * (1 - Tax rate) + Depreciation tax shield

= (3040000 - 507343) * (1 - 25%) + 700000 * 25%

= $2,074,492.75

NPV:

Net Present value = Annual cash flow * PV of $1 annuity for 20 years at 10% rate - Initial investment

= 2074492.75 * (1 - 1 /(1 + 10%)^20) / 10% - 14000000

= $3,661,326.21

Answer 2:

Price of the bond = $1,038.51

Working:

Par value = $1,000

Semiannual coupon = 1000 * 8.11%/2 = $40.55

Number of semiannual period = 17 * 2 = 34

Semiannual yield = 7.70%/ 2 = 3.85%

Price of bond = PV of all coupon payments + PV of redemption value after 34 semiannual periods

= 40.55 * (1 - 1 /(1 + 3.85%)^34)/3.85% + 1000 / (1 + 3.85%)^34

= $1038.51

Answer 3:

Price of the bond = $971.09

Working:

Par value = $1,000

Semiannual coupon = 1000 * 7.81%/2 = $39.05

Number of semiannual period = 10 * 2 = 20

Semiannual yield = 8.24%/ 2 = 4.12%

Price of bond = PV of all coupon payments + PV of redemption amount after 20 semiannual periods

= 39.05 * (1 - 1 /(1 + 4.12%)^20)/4.12% + 1000 / (1 + 4.12%)^20

= $971.09